### v2

@@ 4.4 Comments on JT/Mirzakhani limit
In eberhardt20232d , in particular, an intersection-theoretic formula, inspired by the minimal string answers, as well as “matrix integral” formulation that can be used to compute these volumes were proposed. The results of the present paper suggests that there exists a different approach, in terms of $x-y$ swapped Mirzakhani spectral curve. We consider
$\begin{cases}x^{JT}(z)=\frac{1}{4\pi}\sin 2\pi z\\ y^{JT}(z)=z^{2}\\ \end{cases}$ (63)
and the standard bidifferential; branching points for this curve are the roots of $\cos(2\pi z);z_{m}=\frac{1}{4}(2m+1),m\in\mathbb{Z}$. The local Galois involution ${\overline{z}}(z)$ near $z_{m}$ is simply $2z_{m}-z$. The kernel is then
@@
$-A^{0\,JT}_{3}(\kappa_{1},\kappa_{2},\kappa_{3})=\sum\limits_{m\in\mathbb{Z}}\frac{(-1)^{m}}{2\pi z_{m}}\left(\prod\limits_{i=0}^{2}\cos(2\pi\kappa_{i}z_{m})\right)$ (69)
@@
$\sum\limits_{i=1}^{3}\sum\limits_{\underset{m,n\in\mathbb{Z}}{m\neq n;}}\frac{(-1)^{m+n}}{4\pi^{2}z_{n}z_{m}}\frac{\prod\limits_{j=0,i}\cos(2\pi\kappa_{j}z_{m})\prod\limits_{j\neq 0,i}\cos(2\pi\kappa_{j}z_{n})}{(z_{m}-z_{n})^{2}}-\sum\limits_{m\in\mathbb{Z}}\frac{\left(\prod\limits_{i=0}^{3}\cos(2\pi\kappa_{i}z_{m})\right)}{2z_{m}^{2}}\left(-1+\sum\limits_{i=0}^{3}\kappa_{i}^{2}\right)$ (72)
@@
Let us comment on how one can motivate the formula (58) by explicit calculation in this limit. The general formula in Dunin-Barkowski:2012kbi , representing $\omega_{g,n}$ as sums over stable graphs, says that the factors associated to the edges of the graph are of the form $\sum\limits_{r,s=0}^{\infty}B_{m_{1},2r,m_{2},2s}\psi_{\bullet}^{r}\psi_{\circ}^{s}$, where coefficients $B_{m,r,l,s}$ are determined by expanding the bidifferential $\omega_{0,2}$ in a suitable local coordinate $\zeta$ (it should satisfy ${\overline{\zeta}}^{(m)}(\zeta)=-\zeta$; in our case we can take $\zeta^{(m)}=z-z_{m}$) in the vicinity of branching points:
@@
[…]
@@
$\sum\limits_{n=1}^{\infty}\frac{(-4\pi^{2}\Psi)^{n-1}}{n!}B_{2n}\left(\frac{\kappa}{2}\right)$ (76)
@@ 5.1 Conjecture
Let us first motivate our conjecture. Consider the $n$-point amplitude with the insertion of the simplest ground ring field $O_{1,1}$ — the unit operator (for example, let all other operators be tachyons). Obviously it should be the same as the $(n-1)$-point amplitude, that is calculated by applying (32) to $\check{\omega}_{g,n-1}$. Can we find a linear transformation of the $n$-differential $\check{\omega}_{g,n}$ with respect to $n$-th argument that would produce this result?
@@
[…]
Then, up to a factor that depends on $g$ and $n$ (but does not depend on parameters of other insertions) transformation $\sum\limits_{m}\underset{z=\zeta_{m}}{\text{Res}}\,\Phi_{1}(z)\cdot...$ acts as an insertion of unit operator for any correlator. Note that $\Phi_{1}$ is a polynomial of degree $32p+$ 3; if we extrapolate the correspondence between the degree of the polynomial and the Liouville momentum noted in the footnote after (32), this degree corresponds to unit operator $V_{1,1}$ in the Liouville sector. This is also reassuring.
@@ 5.2 Check against HEM
$\mathcal{N}_{k_{1}\dots k_{5}}^{(0)}\left(\frac{4b^{2}V^{p}_{0,4}(4\pi bP_{1,-k_{2}},\dots 4\pi bP_{1,-k_{5}})}{2\pi^{2}}-2b^{6}(k_{1}^{2}-1)\right)+\frac{2b^{2}}{2}\sum\limits_{i<j,2}^{5}\sum\limits_{n=1}^{2p}G(n)\mathcal{N}^{(0)}_{k_{i}k_{j}n}\mathcal{N}^{(0)}_{n\cdot\cdot\cdot}$ (85)
@@
$\left\langle\mathcal{O}_{1,k_{1}}(x)\int d^{2}z\,\mathcal{U}_{1,k_{2}}(z)\,\mathcal{W}_{a_{1}}(x_{1})\mathcal{W}_{a_{2}}(x_{2})\mathcal{W}_{x_{3}}(\infty)\right\rangle\sim$
@@
Let us compare the HEM answer versus the one that follows from the “conjecture” (85). Consider “generic” values of three tachyon parameters $k_{3},k_{4},k_{5}$ (the ones for which the HEM calculation should apply); “genericness” implies the following for the fusion numbers
@@
$4b^{2}\left(k_{1}\alpha_{0,4}(k_{2}|a_{3},a_{4},a_{5})-\underbrace{4b\sum\limits_{s=|k_{2}-k_{1}|+1:2}^{k_{2}+k_{1}-1}q_{0,s-k_{1}}^{(1,k_{2})}(a_{1,k_{1}})\cdot s}_{(1)}\right)$ (88)
$\alpha_{0,4}$ is
$\frac{V^{p}_{0,4}(4\pi bP_{1,-k_{2}},\dots)}{2\pi^{2}}\mathcal{N}^{(0)}_{k_{2}\dots k_{5}}+\sum\limits_{i=3,4,5}\sum\limits_{l=-k_{2}+1:2}^{k_{2}-1}G(k_{i}+l)$ (89)
(85), on the other hand, reads
$4b^{2}\left(k_{1}\frac{V^{p}_{0,4}(4\pi bP_{1,-k_{2}},\dots)}{2\pi^{2}}+\sum\limits_{i=3,4,5}\sum\limits_{l=-k_{2}+1:2}^{k_{2}-1}\underbrace{k_{1}}_{\mathcal{N}_{k_{1}m\bullet\circ}}G(k_{i}+l)-b^{4}\frac{k_{1}k_{2}(k_{1}^{2}-1)}{2}+\right.$ (90)
@@
$\underbrace{2b\sum\limits_{i=3,4,5}\sum\limits_{l=-(k_{1}-1):2}^{k_{1}-1}\sum\limits_{m=-(k_{2}-1):2}^{k_{2}-1}\left(q_{0,-m}^{(1,k_{2})}(a_{i})-q_{0,-m}^{(1,k_{2})}(a_{i}+lb/2)\right)}_{(2)}$ (93)
@@ 5.3 Discussion
The two answers do not match, although the difference is quite simple. Let us discuss what this might imply if we take the conjecture (85) seriously.
@@
Unlike the HEM expression, the answer (85) is manifestly symmetric under permutations of $k_{2}\dots k_{5}$ and predicts expected discontinuities. An important question then is to resolve the conflict between different answers and compare (85) (as well as HEM prediction) with the integral $\left\langle\mathcal{O}_{1,k_{1}}(x)\int d^{2}z\,\mathcal{U}_{1,k_{2}}(z)\,\mathcal{W}_{a_{1}}(x_{1})\mathcal{W}_{a_{2}}(x_{2})\mathcal{W}_{a_{3}}(x_{3})\right\rangle$ we study.
@@ 6 Conclusions
  • It is of interest to prove the diagrammatic expansion that computes $\check{A}^{g}_{n}$ that is hinted at by obtained expressions, or to find a different way to streamline the calculation of amplitudes in this approach. Perhaps one can adapt the universal $x-y$ swap formula Alexandrov:2022ydc for this purpose.
@@
  • The main observation of this paper raises a question of whether there is an analogous simple way to match correlators in two approaches for the general $(p,q)$ minimal string case. This is a much more subtle issue than for Lee-Yang series: even for three-point amplitudes on the sphere the answer (fusion coefficient) can not be simply reproduced by either the usual or $x-y$ dual spectral curve; instead it is given by the product of the two answers Marshakov:2024lvk . On the other hand, it is very easy to imagine a natural generalization of (61) to this case, by replacing the fusion factors with the appropriate ones for $(p,q)$ model and extending the sum over $k_{e}$ to the lattice of all Liouville momenta of the form $P_{m,n}$. Results of the current paper suggest that ramification points are naturally associated with primary fields in the matter sector; perhaps there is a more convenient TR formulation with $(p-1)(q-1)$ ramification points instead of the standard one coming from the two-matrix model.
@@ Appendix A Cancellation of “on-shell poles” from descendants
×$\times\left(\frac{1}{P_{M}^{2}+b_{\sigma_{B}(1)}}\frac{1}{a_{\sigma_{B}(1)}+b_{\sigma_{B}(2)}}\dots\frac{1}{a_{\sigma_{B}(|B|-1)}+b_{\sigma_{B}(|B|)}}\right)$ (99)
@@ Appendix C Expressions for $\check{A}^{g}_{n}$ for $(g,n)=(0,5)$ and $(1,2)$
×$\times\left(k_{1}(2p+1)+k_{2}\left(2p-k_{2}+1\right)-k_{1}^{2}+p(p+1)+2\right)+$

### v3

@@ 4.4 Comments on JT/Mirzakhani limit
In eberhardt20232d , in particular, an intersection-theoretic formula, inspired by the minimal string answers, as well as “matrix integral” formulation that can be used to compute these volumes were proposed. The results of the present paper suggests that there exists a different approach, in terms of xyx-y swapped Mirzakhani spectral curve. We consider
[…]
and the standard bidifferential; branching points for this curve are the zeroes of $\cos(2\pi z);z_{m}=\frac{1}{4}(2m+1),m\in\mathbb{Z}$. The local Galois involution ${\overline{z}}(z)$ near $z_{m}$ is simply $2z_{m}-z$. The kernel is then
@@
$A^{0\,JT}_{3}(\kappa_{1},\kappa_{2},\kappa_{3})=\sum\limits_{m\in\mathbb{Z}}\frac{(-1)^{m}}{2\pi z_{m}}\left(\prod\limits_{i=0}^{2}\cos(2\pi\kappa_{i}z_{m})\right)$ (69)
@@
$A^{0\,JT}_{4}(\kappa_{0}\dots\kappa_{3})$ $=\sum\limits_{i=1}^{3}\sum\limits_{\underset{m,n\in\mathbb{Z}}{m\neq n;}}\frac{(-1)^{m+n}}{4\pi^{2}z_{n}z_{m}}\frac{\prod\limits_{j=0,i}\cos(2\pi\kappa_{j}z_{m})\prod\limits_{j\neq 0,i}\cos(2\pi\kappa_{j}z_{n})}{(z_{m}-z_{n})^{2}}$-
$-\sum\limits_{m\in\mathbb{Z}}\frac{\left(\prod\limits_{i=0}^{3}\cos(2\pi\kappa_{i}z_{m})\right)}{2z_{m}^{2}}\left(-1+\sum\limits_{i=0}^{3}\kappa_{i}^{2}\right)$ (72)
@@
Let us comment on how one can motivate the formula (58) by explicit calculation in this limit. The general formula in Dunin-Barkowski:2012kbi , representing $\omega_{g,n}$ as sums over stable graphs, says that the factors associated to the edges of the graph are of the form $\sum\limits_{r,s=0}^{\infty}B_{m_{1},2r,m_{2},2s}\psi_{\bullet}^{r}\psi_{\circ}^{s}$, where coefficients $B_{m,r,l,s}$ are determined by expanding the bidifferential $\omega_{0,2}$ in a suitable local coordinate $\zeta$ (it should satisfy ${\overline{\zeta}}^{(m)}(\zeta^{(m)})=-\zeta^{(m)}$; in our case we can take $\zeta^{(m)}=z-z_{m}$) when $z_{1}$ and $z_{2}$ are in the vicinity of branching points $z_{m_{1}}$ and $z_{m_{2}}$:
@@
To get a similar representation to finite $p$ case, with ¡¡fusion coefficients¿¿ entering the vertices, this series should be represented as an integral of the form $\int\limits_{0}^{1}d\kappa\cos(2\pi\kappa z_{m_{1}})\cos(2\pi\kappa z_{m_{2}})\dots$ (this is the “continuous” analogue of $\sum\limits_{k=1}^{p}\sin\pi km_{1}b^{2}\sin\pi km_{2}b^{2}\dots$). This amounts to computing inverse Fourier transform of this series with respect to $(m_{1}-m_{2})$, for which we use
@@
$\sum\limits_{m\in\mathbb{Z}}\frac{\cos(\pi\kappa m)}{\Psi}\sum\limits_{n=1}^{\infty}\frac{(2n)!}{n!}\frac{\Psi^{n}}{m^{2n}}\delta_{m\neq 0}=\sum\limits_{n=1}^{\infty}\frac{(-4\pi^{2}\Psi)^{n-1}}{n!}B_{2n}\left(\frac{\kappa}{2}\right)$ (76)
@@ 5.1 Conjecture
Let us first motivate our conjecture. Consider the $n$-point amplitude with the insertion of the simplest ground ring field $O_{1,1}$ — the unit operator (for example, let all other operators be tachyons). Obviously it should be the same as the (n1)(n-1)-point amplitude, that is calculated by applying (32) to $\check{\omega}_{g,n-1}$. Can we find a linear transformation of the $n$-differential $\check{\omega}_{g,n}$ with respect to $n$-th argument that would produce this result?
@@
where $\Phi=\int\check{y}\,d\check{x}$. Using trigonometric form of Chebyshev polynomials
$\Phi_{1}(z=\cos\theta)=-(2p+1)\int\cos(2\theta)\sin((2p+1)\theta)d\theta=\frac{2p+1}{2}\left(\frac{T_{2p+3}(z)}{2p+3}+\frac{T_{2p-1}(z)}{2p-1}\right)$ (78)
Then, up to a factor that depends on $g$ and $n$ (but does not depend on parameters of other insertions) transformation $\sum\limits_{m}\underset{z=\zeta_{m}}{\text{Res}}\,\Phi_{1}(z)\cdot$... acts as an insertion of unit operator for any correlator. Note that $\Phi_{1}$ is a polynomial of degree $2p+3$; if we extrapolate the correspondence between the degree of the polynomial and the Liouville momentum noted in the footnote after (32), this degree corresponds to unit operator $V_{1,1}$ in the Liouville sector. This is also reassuring.
@@ 5.2 Check against HEM
$\frac{1}{4b^{4}}\text{Res}\left(\Phi_{k_{1}}(z_{1})\left(\prod\limits_{i=2}^{5}\frac{T_{2(p-k_{i})+1}(z_{i})}{2(p-k_{i})+1}\right)\check{\omega}_{0,5}(z_{1}\dots z_{5})\right)=$
$=\mathcal{N}_{k_{1}\dots k_{5}}^{(0)}\left(\frac{4b^{2}V^{p}_{0,4}(4\pi bP_{1,-k_{2}},\dots 4\pi bP_{1,-k_{5}})}{2\pi^{2}}-2b^{6}(k_{1}^{2}-1)\right)+\frac{2b^{2}}{2}\sum\limits_{i<j,2}^{5}\sum\limits_{n=1}^{2p}G(n)\mathcal{N}^{(0)}_{k_{i}k_{j}n}\mathcal{N}^{(0)}_{n\cdot\cdot\cdot}$ (85)
@@
$\left\langle\mathcal{O}_{1,k_{1}}(x)\int d^{2}z\,\mathcal{U}_{1,k_{2}}(z)\,\mathcal{W}_{a_{1}}(x_{1})\mathcal{W}_{a_{2}}(x_{2})\mathcal{W}_{x_{3}}(\infty)\right\rangle_{\text{HEM}}\sim$
@@
Let us compare the HEM answer versus the one that follows from the “conjecture” (5.2). Consider “generic” values of three tachyon parameters $k_{3},k_{4},k_{5}$ (the ones for which the HEM calculation should apply); “genericness” implies the following for the fusion numbers
@@
$\left\langle\mathcal{O}_{1,k_{1}}(x)\int d^{2}z\,\mathcal{U}_{1,k_{2}}(z)\,\mathcal{W}_{a_{1}}(x_{1})\mathcal{W}_{a_{2}}(x_{2})\mathcal{W}_{x_{3}}(\infty)\right\rangle_{\text{HEM}}=$
$=4b^{2}\left(k_{1}\alpha_{0,4}(k_{2}|a_{3},a_{4},a_{5})-\underbrace{4b\sum\limits_{s=|k_{2}-k_{1}|+1:2}^{k_{2}+k_{1}-1}q_{0,s-k_{1}}^{(1,k_{2})}(a_{1,k_{1}})\cdot s}_{(1)}\right)$ (88)
where
$\alpha_{0,4}(k_{2}|a_{3},a_{4},a_{5})=\frac{V^{p}_{0,4}(4\pi bP_{1,-k_{2}},\dots)}{2\pi^{2}}\mathcal{N}^{(0)}_{k_{2}\dots k_{5}}+\sum\limits_{i=3,4,5}\sum\limits_{l=-k_{2}+1:2}^{k_{2}-1}G(k_{i}+l)$ (89)
(5.2), on the other hand, reads
$\frac{1}{4b^{4}}\text{Res}\left(\Phi_{k_{1}}(z_{1})\left(\prod\limits_{i=2}^{5}\frac{T_{2(p-k_{i})+1}(z_{i})}{2(p-k_{i})+1}\right)\check{\omega}_{0,5}(z_{1}\dots z_{5})\right)=$
$=4b^{2}\left(k_{1}\frac{V^{p}_{0,4}(4\pi bP_{1,-k_{2}},\dots)}{2\pi^{2}}+\sum\limits_{i=3,4,5}\sum\limits_{l=-k_{2}+1:2}^{k_{2}-1}\underbrace{k_{1}}_{\mathcal{N}_{k_{1}m\bullet\circ}}G(k_{i}+l)-b^{4}\frac{k_{1}k_{2}(k_{1}^{2}-1)}{2}+\right$. (90)
@@
$-b^{2}\sum\limits_{i=3,4,5}\sum\limits_{l=-k_{2}+1:2}^{k_{2}-1}\sum\limits_{m=-k_{1}+1:2}^{k_{1}-1}\left(|p+\frac{1}{2}-k_{i}-m-l|-|p+\frac{1}{2}-k_{i}-l|\right)=$
$=\underbrace{2b\sum\limits_{i=3,4,5}\sum\limits_{l=-(k_{1}-1):2}^{k_{1}-1}\sum\limits_{m=-(k_{2}-1):2}^{k_{2}-1}\left(q_{0,-m}^{(1,k_{2})}(a_{i})-q_{0,-m}^{(1,k_{2})}(a_{i}+lb/2)\right)}_{(2)}$ (93)
@@ 5.3 Discussion
The two answers do not match, although the difference is quite simple. Let us discuss what this might imply if we take the conjecture (5.2) seriously.
@@
Unlike the HEM expression, the answer (5.2) is manifestly symmetric under permutations of $k_{2}\dots k_{5}$ and predicts expected discontinuities. An important question then is to resolve the conflict between different answers and compare (5.2) (as well as HEM prediction) with the integral $\left\langle\mathcal{O}_{1,k_{1}}(x)\int d^{2}z\,\mathcal{U}_{1,k_{2}}(z)\,\mathcal{W}_{a_{1}}(x_{1})\mathcal{W}_{a_{2}}(x_{2})\mathcal{W}_{a_{3}}(x_{3})\right\rangle$ we study.
@@ 6 Conclusions
  • It is of interest to prove the diagrammatic expansion that computes $\check{A}^{g}_{n}$ that is hinted at by obtained expressions, or to find a different way to streamline the calculation of amplitudes in this approach. Perhaps one can adapt the universal xyx-y swap formula Alexandrov:2022ydc for this purpose.
@@
  • The main observation of this paper raises a question of whether there is an analogous simple way to match correlators in two approaches for the general $(p,q)$ minimal string case. This is a much more subtle issue than for Lee-Yang series: even for three-point amplitudes on the sphere the answer (fusion coefficient) can not be simply reproduced by either the usual or xyx-y dual spectral curve; instead it is given by the product of the two answers Marshakov:2024lvk . On the other hand, it is very easy to imagine a natural generalization of (61) to this case, by replacing the fusion factors with the appropriate ones for $(p,q)$ model and extending the sum over $k_{e}$ to the lattice of all Liouville momenta of the form $P_{m,n}$. Results of the current paper suggest that ramification points are naturally associated with primary fields in the matter sector; perhaps there is a more convenient TR formulation with (p1)(q1)(p-1)(q-1) ramification points instead of the standard one coming from the two-matrix model.
@@ Appendix A Cancellation of “on-shell poles” from descendants
$\times\left(\frac{1}{P_{M}^{2}+b_{\sigma_{B}(1)}}\frac{1}{a_{\sigma_{B}(1)}+b_{\sigma_{B}(2)}}\dots\frac{1}{a_{\sigma_{B}(|B|-1)}+b_{\sigma_{B}(|B|)}}\right)$ (99)
@@ Appendix C Expressions for $\check{A}^{g}_{n}$ for $(g,n)=(0,5)$ and $(1,2)$
$\times\left(k_{1}(2p+1)+k_{2}\left(2p-k_{2}+1\right)-k_{1}^{2}+p(p+1)+2\right)+$